### v1

# Picosecond Wireless Synchronization with Entangled Photons via Grid-Based Quantum Coverage in Indoor Optical Systems

Abstract
In this paper, we propose a novel entanglement-assisted synchronization framework for indoor optical wireless systems based on grid-based beam steering. A central transmitter equipped with a spontaneous parametric down-conversion (SPDC) source emits time-energy entangled photon pairs, directing user photons toward spatially defined grids based on estimated user positions, while retaining reference photons for timestamping. The room is partitioned into multiple beam-aligned regions, and photon reception probability is analytically modeled using local coordinate transformations and Gaussian beam optics. To address the inherent sparsity and randomness of photon detection, we develop a robust two-stage synchronization algorithm that performs sparse bit-pattern matching and timestamp averaging to estimate timing offsets. Monte Carlo simulations are then performed to evaluate the impact of synchronization duration, grid resolution, and photon-pair generation rate on synchronization accuracy. Results show that finer grid configurations and optimal photon-pair rates significantly reduce synchronization error, achieving sub-$10$ ps timing accuracy within $1\,\text{ms}$ of synchronization time. The proposed approach enables scalable, high-precision quantum synchronization in wireless indoor environments, laying the groundwork for future joint quantum communication, sensing, and positioning systems.
Index Terms:
Entangled photons, grid of beams, indoor optical wireless systems, quantum time transfer, quantum synchronization, single-photon detection, SPDC source.
I Introduction
The convergence of quantum information science and wireless communication is driving the development of a new class of systems that promise secure communication, high-precision sensing, and distributed quantum computing [1, 2]. Among these, optical wireless communication (OWC) systems, which use light for high-speed data transfer, have emerged as a promising platform for enabling quantum-enhanced applications, including quantum key distribution (QKD), quantum sensing, and local-area quantum networks [2, 3, 4, 5, 6]. A fundamental requirement across nearly all such applications is the ability to distribute and detect entangled or single photons with extremely high temporal precision, often at the picosecond scale [7].
This stringent requirement for temporal precision becomes especially critical in protocols such as QKD and entanglement swapping, where synchronization between a central node and multiple users must reach the picosecond level. In these scenarios, the identification of correlated photon detection events depends on extremely narrow coincidence windows. Even minor clock offsets or uncertainties in propagation delay can significantly degrade system performance, increasing error rates and undermining quantum coherence [8, 9]. As a result, time synchronization is not merely an engineering challenge but a fundamental physical necessity in quantum communication systems.
Given the critical role of synchronization in quantum systems, the field of quantum clock synchronization (QCS) has evolved along two primary paths. The first focuses on long-haul fiber-optic quantum links, where weak quantum signals must coexist with classical data traffic [10, 11]. These systems employ techniques such as wavelength division multiplexing and polarization management to mitigate crosstalk and dispersion effects. Recent advances have further extended this, where quantum-enhanced two-way time transfer leverages the strong temporal correlations in energy-time entangled photons to achieve precise synchronization stability [12, 13]. The second path targets long-distance free-space quantum links, including ground-to-satellite communication, where atmospheric turbulence, solar background radiation, and Doppler shifts pose significant challenges. Solutions in this domain typically involve adaptive optics, narrowband filtering, time transfer, and probabilistic error correction [14, 15, 16]. While both approaches are technically sophisticated, their assumptions and mitigation strategies are highly specialized and not directly applicable to indoor environments.
While existing QCS approaches have focused on long-haul fiber or free-space links, a parallel line of development in the classical domain has explored scalable indoor OWC architectures. Among these, the Grid-of-Beams (GoB) approach has emerged as a practical solution for supporting mobile users [17, 18, 19, 20, 21]. In a GoB system, a dense array of static, narrow laser beams provides full spatial coverage, and user mobility is managed through digital handovers between adjacent grid cells, thereby eliminating the need for continuous beam tracking [17, 19]. However, these systems are inherently quantum-agnostic. Each handover alters the optical path length, disrupting prior timing calibration. Moreover, current GoB architectures lack mechanisms to rapidly re-establish the picosecond-level synchronization required by quantum protocols. More broadly, existing QCS schemes assume stable links and long integration times, making them unsuitable for dynamic indoor environments characterized by frequent handovers and ambient lighting noise. This highlights a critical gap at the intersection of quantum synchronization and mobile indoor OWC that must be addressed to enable practical quantum networking in real-world settings.
To bridge this gap, we propose the first integrated framework for entanglement-assisted picosecond synchronization tailored to mobile indoor optical wireless environments. Our approach leverages the temporal correlations of entangled photon pairs in conjunction with a scalable optical delivery system to enable precise synchronization. This method allows for accurate estimation of timing parameters without relying on a shared clock or fixed spatial configuration, making it well-suited for dynamic, mobile scenarios.
To enable practical deployment in dynamic indoor environments, we build upon existing and experimentally validated GoB architectures, originally developed for classical OWC systems to provide full indoor coverage and support user mobility [17, 22]. The GoB discretizes space into fixed beams, each targeting a predefined coverage cell, thereby simplifying link acquisition and handover. We extend this architecture by embedding a quantum-native synchronization layer, enabling support for entanglement-based protocols. We complement this architecture with a two-stage synchronization algorithm designed to operate over sparse and noisy single-photon detection streams. The first stage performs coarse alignment using bit-pattern matching, while the second refines synchronization through timestamp averaging. To support system design and performance evaluation, we also develop a comprehensive analytical model that captures photon-pair generation statistics, Gaussian beam propagation, user position uncertainty, and timing jitter from both the source and detectors.
To the best of our knowledge, this is the first work to present and analyze in detail an entanglement-assisted synchronization framework tailored to a grid-based indoor OWC system. Our main contributions are as follows:
  • We propose a complete system architecture for high-precision timing in indoor environments, integrating a grid-based optical delivery system with an entanglement-based synchronization protocol. This design directly addresses the challenges of user mobility and line-of-sight link maintenance.
  • We develop a detailed, end-to-end analytical model that captures the key physical phenomena. This includes the Poisson statistics of the quantum unit, Gaussian beam optics for photon propagation, the impact of user positioning errors on reception probability, and a rigorous treatment of both source and detector timing jitter.
  • We design a practical algorithm specifically to handle the sparse and noisy data characteristic of single-photon channels. The algorithm performs sparse bit-pattern matching for coarse slot-level alignment, followed by timestamp averaging over high-confidence matched events to achieve fine-grained, picosecond-level precision.
  • Through Monte Carlo simulations, we evaluate the performance of our framework under realistic conditions. We demonstrate that sub-$10$ picosecond synchronization accuracy is achievable within a $1$ ms timeframe. Our results quantify the significant impact of spatial grid resolution and reveal the existence of an optimal photon-pair generation rate that balances signal availability against multi-photon ambiguity.
Thus, this work provides a foundational step toward building practical, scalable, and high-precision quantum networks in local indoor settings, bridging quantum timing principles with proven classical networking architectures.
The remainder of this paper is structured as follows. Section II presents the system model considered in this study. Section III describes the spatial modelling of photon reception in the communication setup. Section IV introduces the proposed synchronization models. Simulation results are discussed in Section V. Finally, Section VI concludes the paper and outlines directions for future research.
TABLE I: List of Key Notations
Physical Environment
$x,y,z$ Global coordinate system
$x^{\prime},y^{\prime},z^{\prime}$ Beam-aligned local coordinate system
$L,W,H$ Room length, width, and height
$h_{\text{cov}}$ Optical coverage height from the floor
$N_{x},N_{y}$ Number of grids along $x$ and $y$ axes
$N_{g}=N_{x}\times N_{y}$ Total number of spatial grid zones
$\mathbf{p}_{\text{tx}}$ Transmitter location (origin)
$\mathbf{p}_{u}$ True user position in global coordinates
$\hat{\mathbf{p}}_{u}$ Estimated user position
$\sigma_{p}^{2}$ Position estimation error variance
$\mathbf{d}^{(i,j)}$ Beam direction vector toward grid $(i,j)$
$\mathbf{R}^{(i,j)}$ Rotation matrix to beam-aligned coordinates
$\mathbf{p}_{u}^{\prime}$ User position in local (beam) frame
$r_{a}$ Receiver aperture radius
Transmitter Parameters
$\mu_{t}$ Mean photon-pair generation rate per slot
$\tau_{p}$ Pump pulse width
$t_{g}[k]$ Photon-pair generation time in slot $k$
$\sigma_{t}$ SPDC source temporal jitter
$t_{r}[k]$ Reference photon detection time at transmitter
$n_{r}[k]$ Jitter at transmitter SPAD (reference)
$\sigma_{d}$ SPAD detector timing jitter
$\lambda$ Photon wavelength
$w_{0}$ Beam waist at source
$w(z)$ Beam width at distance $z$
Receiver Parameters
$P_{\text{recept}}^{(i,j)}$ Photon reception probability in grid $(i,j)$
$\eta_{d}$ SPAD quantum detection efficiency
$\mu_{s}^{(i,j)}$ Mean number of detected signal photons per slot
$\mu_{b}$ Mean number of background photons per slot
$N_{s}[k]$ Number of received signal photons in slot $k$
$N_{b}[k]$ Number of background photons in slot $k$
$t_{u}[k]$ Photon detection timestamp at user
$n_{u}[k]$ Receiver SPAD timing jitter
$\tau_{\text{sync}}$ Effective synchronization offset
II System Model
We consider an indoor quantum synchronization system based on a grid-structured optical wireless delivery architecture that utilizes time-energy entangled photon pairs. A central quantum unit, mounted on the ceiling, contains a spontaneous parametric down-conversion (SPDC) source that continuously generates entangled photons. In each pair, one photon remains at the central unit as a precise timing reference, while the other is directed toward the user through one of $N_{g}$ spatially distinct optical wireless transmission paths. Each transmission path corresponds to a fixed illumination region (grid) on the indoor surface, effectively dividing the room into $N_{g}$ non-overlapping zones. In the following subsections, we provide a detailed description of the system model and the parameters considered in this paper. Meanwhile, a summary of the notations used is presented in Table I.
Figure 1: System architecture of the proposed entanglement-based indoor quantum synchronization scheme. A central ceiling-mounted transmitter emits time-energy entangled photon pairs, directing the user photon toward a spatially selected beam corresponding to the estimated user location. The beam is aligned with a specific grid region in the room, while the reference photon is time-stamped at the transmitter. The receiver structure includes a circular aperture and SPAD array aligned to detect the photon and record its arrival time for synchronization. Local $(x^{\prime},y^{\prime},z^{\prime})$ and global $(x,y,z)$ coordinate systems are depicted for spatial modeling.
II-A Geometric Configuration and Coordinate Framework
We consider a rectangular indoor environment with length $L$, width $W$, and height $H$. The quantum transmitter unit is mounted at the center of the ceiling, which is chosen as the origin of the global 3D coordinate system. Throughout this work, we denote global positions using the coordinate system $(x,y,z)$, with the transmitter located at $\mathbf{p}_{\text{tx}}=[0,0,0]$. The room is partitioned into $N_{g}=N_{x}\times N_{y}$ non-overlapping grid zones, where $N_{x}$ and $N_{y}$ are the number of divisions along the $x$ and $y$ axes, respectively. Each grid is centered at a position $\mathbf{p}_{g}^{(i,j)}$ on the horizontal coverage plane $z=-H+h_{\text{cov}}$, where $1\leq i\leq N_{x}$ and $1\leq j\leq N_{y}$. Thus, we have
$\mathbf{p}_{g}^{(i,j)}=\left(\left(i-\tfrac{1}{2}\right)\Delta x-\tfrac{L}{2},\;\left(j-\tfrac{1}{2}\right)\Delta y-\tfrac{W}{2},\;-H+h_{\text{cov}}\right),$ (1)
with $\Delta x=\frac{L}{N_{x}}$ and $\Delta y=\frac{W}{N_{y}}$ being the width and height of each grid zone. Here, $h_{\text{cov}}$ denotes the vertical distance above the floor at which optical coverage is optimized, i.e., the height where the beam footprint is best aligned with the grid layout. This coverage design follows the 3D optical model and beam divergence analysis discussed in [21].
Each optical wireless transmission path is aligned to direct a photon beam from $\mathbf{p}_{\text{tx}}$ to the center of a corresponding grid. The direction vector of the beam toward grid $(i,j)$ is given by
$\mathbf{d}^{(i,j)}=\mathbf{p}_{g}^{(i,j)}-\mathbf{p}_{\text{tx}}=\mathbf{p}_{g}^{(i,j)}.$ (2)
To ensure that the entire grid at $h_{\text{cov}}$ is covered by the optical beam, the divergence angle $\phi_{\text{div}}^{(i,j)}$ of each beam is adjusted based on its propagation vector $\mathbf{d}^{(i,j)}$. The divergence angle of each beam is tuned to ensure complete coverage at the receiver plane $z=h_{\text{cov}}$. For further geometric considerations and coverage performance details, refer to [21].
Let $\mathbf{p}_{u}=[x_{u},y_{u},z_{u}]$ denote the true position of the user, and let $\hat{\mathbf{p}}_{u}=[\hat{x}_{u},\hat{y}_{u},\hat{z}_{u}]$ represent the estimated position available to the system. The estimation is subject to random errors modeled as independent Gaussian noise in each coordinate. Specifically, we have
$\hat{x}_{u}=x_{u}+n_{x},\quad\hat{y}_{u}=y_{u}+n_{y},\quad\hat{z}_{u}=z_{u}+n_{z},$ (3)
where $n_{x},n_{y},n_{z}\sim\mathcal{N}(0,\sigma_{p}^{2})$ are zero-mean Gaussian random variables with identical variance $\sigma_{p}^{2}$. These represent the initial spatial uncertainty in user localization arising from sensing or tracking limitations. Given $\hat{\mathbf{p}}_{u}$, the system activates the optical transmission path corresponding to the grid $(i^{\star},j^{\star})$ that minimizes the angular deviation between the beam direction $\mathbf{d}^{(i,j)}$ and the estimated user vector. Formally, the selected grid is given by
$(i^{\star},j^{\star})=\arg\min_{(i,j)}\;\theta_{\text{err}}^{(i,j)}=\arg\min_{(i,j)}\;\cos^{-1}\left(\frac{\mathbf{d}^{(i,j)}\cdot\hat{\mathbf{d}}_{u}}{\|\mathbf{d}^{(i,j)}\|\|\hat{\mathbf{d}}_{u}\|}\right),$ (4)
where $\hat{\mathbf{d}}_{u}=\hat{\mathbf{p}}_{u}-\mathbf{p}_{\text{tx}}$ is the direction vector from the transmitter to the estimated user location. This selection ensures that the beam is directed toward the grid whose nominal axis most closely aligns with the estimated user position in 3D space.
Moreover, to simplify the analysis of photon propagation, we transform all global positions from the original $(x,y,z)$ coordinate system to a local coordinate frame $(x^{\prime},y^{\prime},z^{\prime})$ aligned with the beam direction $\mathbf{d}^{(i,j)}$. In this frame, the $z^{\prime}$-axis is parallel to $\mathbf{d}^{(i,j)}$, and the $x^{\prime}$-$y^{\prime}$ plane is perpendicular to it. The user position is then transformed from the global coordinate system to the beam-aligned local frame using
$\mathbf{p}_{u}^{\prime}=\mathbf{R}^{(i,j)}\left(\mathbf{p}_{u}-\mathbf{p}_{\text{tx}}\right),$ (5)
where $\mathbf{p}_{u}^{\prime}=[x_{u}^{\prime},y_{u}^{\prime},z_{u}^{\prime}]^{\top}$ represents the user coordinates in the local $(x^{\prime},y^{\prime},z^{\prime})$ frame, where $\cdot$$^{\top}$ denotes the transpose operator. The construction of the rotation matrix $\mathbf{R}^{(i,j)}$ is provided in Appendix A.
II-B Transmitter Architecture and Photon Pair Emission
The central transmitter unit employs a SPDC source to generate time-energy entangled photon pairs. Each generated pair consists of two photons: one designated as the reference photon and the other as the user photon. The reference photon is routed to a time-stamping detection module co-located at the transmitter, while the user photon is directed toward a spatially-resolved optical output based on the user’s estimated position, as shown in Fig. 1.
A SPDC source is widely adopted for quantum communication and synchronization due to its scalability and compatibility with free-space optics [23]. The system is driven by a periodic train of pump pulses, each with temporal width $\tau_{p}$. We model the photon-pair generation time $t_{g}$ as a Gaussian random variable centered at the pump pulse peak $t_{p}$, with standard deviation $\sigma_{t}$, as follows [24]
$t_{g}\sim\mathcal{N}(t_{p},\sigma_{t}^{2}).$ (6)
This temporal uncertainty arises from the finite spectral bandwidth of the pump and the phase-matching conditions in the nonlinear crystal [25]. The jitter $\sigma_{t}$ represents an intrinsic timing error that limits the accuracy of photon-level synchronization. Typical values of $\sigma_{t}$ range from tens to hundreds of femtoseconds, depending on the pump-cavity configuration and spectral filtering [26]. To ensure slot-level temporal isolation and mitigate timing ambiguity, each quantum slot is defined to be longer than the pump pulse duration, i.e., $T_{qs}>\tau_{p}$. Specifically, the slot duration is modeled as
$T_{qs}=t_{\text{guard}}^{(1)}+\tau_{p}+t_{\text{guard}}^{(2)},$ (7)
where $t_{\text{guard}}^{(1)}$ and $t_{\text{guard}}^{(2)}$ denote pre- and post-pulse guard intervals, respectively. These intervals are selected to absorb both the temporal jitter $\sigma_{t}$ of the SPDC source and the detection-time uncertainty $\sigma_{d}$ of the reference detector (e.g., SPAD).
Given the spontaneous nature of SPDC, the number of photon pairs generated within each quantum slot is random. Specifically, the emission statistics follow a Poisson distribution [27]
$\mathbb{P}(N_{t}=k)=\frac{\mu_{t}^{k}}{k!}e^{-\mu_{t}},\quad\text{where }\mu_{t}<1.$ (8)
Here, $\mu_{t}$ denotes the mean number of photon pairs per slot.
Furthermore, each photon in the pair encodes a quantum bit (qubit) $|\psi\rangle$ in the polarization basis, defined as
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,\quad\text{with }|\alpha|^{2}+|\beta|^{2}=1,$ (9)
where $|0\rangle$ and $|1\rangle$ represent the horizontal and vertical polarization states, respectively. The entangled state of the pair can be written as:
$|\Psi^{-}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{r}|1\rangle_{u}-|1\rangle_{r}|0\rangle_{u}\right)$ (10)
where subscripts $r$ and $u$ denote the reference and user photons, respectively.
To direct the user photon toward the appropriate receiver location, the transmitter utilizes a dynamic beam selection mechanism based on estimated user coordinates $\hat{\mathbf{p}}_{u}$. As described in the previous subsection, the system activates the optical beam corresponding to the selected grid, minimizing the angular deviation between the estimated user direction and the grid axis (see (4)).
The switching mechanism at the transmitter steers the user photon to the beam path $\mathbf{d}^{(i^{\star},j^{\star})}$ via a high-speed optical switch array. Let $\mathcal{S}_{ij}$ denote the switch corresponding to path $(i,j)$. At each quantum slot, only one $\mathcal{S}_{ij}$ is activated. Simultaneously, the reference photon is routed internally to a time-reference block, where it is detected by a high-resolution single-photon detector (e.g., SPAD). The detection time $t_{r}$ is recorded and serves as a synchronization anchor for the corresponding user photon, allowing for precise time-of-arrival (ToA) estimation upon reception.
II-C Receiver Architecture and Photon Capture
The user-side receiver, shown on the left-hand side of Fig. 1, is designed to detect entangled photons transmitted via spatially directed beams from the central transmitter. The core component of the receiver is a circular aperture that captures incident photons based on their spatial alignment with the selected transmission beam. Each photon emitted toward the user is received with a probability $P_{\text{recept}}^{(i,j)}$, which depends on the alignment between the beam path $(i,j)$ and the user’s actual position. The analytical model for $P_{\text{recept}}^{(i,j)}$ is developed in the following section.
After entering the aperture, the photon propagates through a passive optical path that directs it toward a detection module comprising an array of SPADs. These detectors are polarization-resolving and are used to determine the quantum state of the incoming photon based on its polarization encoding [28]. Each SPAD channel corresponds to a measurement basis (e.g., horizontal or vertical), allowing the receiver to recover the transmitted qubit from the spatially separated entangled photon pair.
Beyond merely detecting the presence of a photon, SPADs employed in this system are capable of precise time-tagging. Upon photon impact, an avalanche current is triggered within the diode, and the timestamp $t_{d}$ corresponding to the avalanche peak (or threshold crossing) is recorded. In practice, SPADs exhibit inherent temporal jitter due to both carrier avalanche buildup and electronic readout delays. This timing uncertainty is typically modeled as a Gaussian random variable $\sigma_{d}$ [29], often referred to as the detector timing jitter
$t_{d}=t_{a}+n_{d},\quad n_{d}\sim\mathcal{N}(0,\sigma_{d}^{2}),$ (11)
where $t_{a}$ is the actual photon arrival time, and $t_{d}$ is the recorded detection time. The value of the standard deviation parameter, $\sigma_{d}$, depends on the SPAD’s material and circuit design, and typically ranges from $20$ ps to $100$ ps in modern systems [30, 31].
III Spatial Modeling of Photon Reception via Aperture
The photon beam emitted by the transmitter is directed toward the center of the selected grid $(i,j)$, following the predefined transmission vector $\mathbf{d}^{(i,j)}$ in (2). Although the photon nominally propagates along $\mathbf{d}^{(i,j)}$, quantum diffraction causes its presence to be distributed laterally in the $x^{\prime}$-$y^{\prime}$ plane, with the probability density decaying radially from the beam axis [32]. This spatial uncertainty is captured by modeling the photon’s transverse wavefunction as a Gaussian mode [33]. Specifically, the spatial probability density in the orthogonal plane at axial distance $z_{u}^{\prime}$ is given by
$P_{\text{spatial}}^{(i,j)}(x^{\prime},y^{\prime})=\frac{2}{\pi w^{2}(z_{u}^{\prime})}\exp\left(-\frac{2(x^{\prime 2}+y^{\prime 2})}{w^{2}(z_{u}^{\prime})}\right),$ (12)
where $w(z_{u}^{\prime})$ denotes the beam width at distance $z_{u}^{\prime}$ along the propagation axis and is expressed as
$w(z_{u}^{\prime})=w_{0}\sqrt{1+\left(\frac{z_{u}^{\prime}\lambda}{\pi w_{0}^{2}}\right)^{2}},$ (13)
with $w_{0}$ being the beam waist (minimum spot size at the focus), and $\lambda$ the photon wavelength.
The receiver aperture, centered at the user’s position $\mathbf{p}_{u}^{\prime}=[x_{u}^{\prime},y_{u}^{\prime},z_{u}^{\prime}]^{\top}$ in the beam-aligned coordinate frame, is assumed to lie in the plane orthogonal to the beam axis, i.e., the $x^{\prime}$-$y^{\prime}$ plane. The aperture captures photons based on their spatial distribution relative to the beam center. Given the radial symmetry of the beam in the $x^{\prime}$-$y^{\prime}$ plane, the photon detection probability depends on the radial displacement of the user from the beam center. This displacement is quantified by the lateral components $x_{u}^{\prime}$ and $y_{u}^{\prime}$ of the user’s position in the local frame.
Let $\rho_{u}^{\prime}=\sqrt{(x_{u}^{\prime})^{2}+(y_{u}^{\prime})^{2}}$ denote the radial offset of the receiver center from the beam axis. The photon detection probability is obtained by integrating the spatial density function (12) over the receiver aperture, which is modeled as a circular disk of radius $r_{a}$ centered at $(x_{u}^{\prime},y_{u}^{\prime})$ in the $x^{\prime}$-$y^{\prime}$ plane. The detection probability is thus given by
$P_{\text{recept}}^{(i,j)}=\iint_{\mathcal{A}_{u}}P_{\text{spatial}}^{(i,j)}(x^{\prime},y^{\prime})\,dx^{\prime}\,dy^{\prime},$ (14)
where $\mathcal{A}_{u}$ denotes the circular region defined by
$\mathcal{A}_{u}=\left\{(x^{\prime},y^{\prime})\in\mathbb{R}^{2}:(x^{\prime}-x_{u}^{\prime})^{2}+(y^{\prime}-y_{u}^{\prime})^{2}\leq r_{a}^{2}\right\}.$ (15)
When the beam width at the receiver plane is much larger than the aperture size, i.e., $w(z_{u}^{\prime})>4r_{a}$, the beam’s intensity doesn’t change much across the tiny area of the aperture [34]. Thus, the spatial variation of the probability density across the receiver area becomes negligible. In this regime, the integral in (14) can be approximated as
$P_{\text{recept}}^{(i,j)}\approx P_{\text{spatial}}^{(i,j)}(x_{u}^{\prime},y_{u}^{\prime})\cdot\pi r_{a}^{2},$ (16)
where $P_{\text{spatial}}^{(i,j)}(x_{u}^{\prime},y_{u}^{\prime})$ is evaluated using (12).
IV Synchronization Model
In order to perform entanglement-based synchronization between the transmitter and the user receiver, we first define a discrete time frame consisting of a sequence of $N_{s}$ quantum time slots. Each slot is intended to carry a potential photon-pair emission by the source, with time-stamping performed by the reference unit located at the transmitter side.
IV-A Slot-Based Reference Sequence Construction
Let $\mathcal{S}=\{1,2,\dots,N_{s}\}$ denote the set of all quantum slots, and for each slot $k\in\mathcal{S}$, let $N_{t}[k]$ denote the number of photon pairs generated. As described earlier in (8), this number follows a Poisson distribution with mean $\mu_{t}<1$. For synchronization purposes, the reference unit maintains two sequences as follows:
  • A data detection sequence $\mathbf{d}_{r}=[d_{r}[1],d_{r}[2],\dots,d_{r}[N_{s}]]$
  • A timestamp sequence $\mathbf{t}_{r}=[t_{r}[1],t_{r}[2],\dots,t_{r}[N_{s}]]$
where $d_{r}[k]\in\{0,1,\texttt{fail}\}$ indicates the detection status of a valid reference photon in slot $k$, and $t_{r}[k]$ is the recorded timestamp when $d_{r}[k]=1$. A quantum time slot $k$ is marked as fail under the two following conditions:
  • No photon pair was generated ($N_{t}[k]=0$)
  • More than one photon pair was generated ($N_{t}[k]\geq 2$)
Accordingly, only those slots for which $N_{t}[k]=1$ are considered valid and eligible for synchronization. The rest are marked as unreliable. Therefore, we have
[…]
Due to physical limitations of the SPAD detector at the reference unit, the timestamp $t_{r}[k]$ is not the exact emission time but includes additive Gaussian noise. Specifically, for each slot $k$ with $d_{r}[k]=1$, the recorded time is
$t_{r}[k]=t_{g}[k]+n_{r}[k],\quad n_{r}[k]\sim\mathcal{N}(0,\sigma_{d}^{2})$ (18)
where $t_{g}[k]$ is the true photon generation time for slot $k$, and $n_{r}[k]$ models the detection-time jitter at the reference SPAD. It is important to note that the generation time $t_{g}[k]$ itself is also a random variable due to intrinsic temporal uncertainty of the SPDC process, modeled in (6). Thus, we have
$t_{g}[k]\sim\mathcal{N}(t_{p}[k],\sigma_{t}^{2})$ (19)
where $t_{p}[k]$ is the center of the pump pulse for slot $k$. Consequently, the effective timestamp $t_{r}[k]$ observed at the reference is subject to a compound noise
$t_{r}[k]\sim\mathcal{N}(t_{p}[k],\sigma_{t}^{2}+\sigma_{d}^{2})$ (20)
which combines the SPDC generation jitter and the SPAD detection jitter. In contrast, the detection noise at the user’s SPAD is assumed independent and modeled separately.
IV-B User-Side Detection Sequence
In each quantum time slot $k$, the number of signal photons $N_{s}[k]$ arriving at the aperture of the user receiver follows a Poisson distribution. Its probability mass function is given by
$\mathbb{P}(N_{s}[k]=n)=\frac{(\mu_{s}^{(i,j)})^{n}}{n!}e^{-\mu_{s}^{(i,j)}},\quad n=0,1,2,\dots$ (21)
where the mean value $\mu_{s}^{(i,j)}$ depends on three factors: the mean photon-pair generation rate $\mu_{t}$ at the transmitter (see (8)), the photon reception probability $P_{\text{recept}}^{(i,j)}$ determined by spatial alignment (see (16)), and the photon detection efficiency $\eta_{d}$ of the user’s SPAD detector. Specifically,
$\mu_{s}^{(i,j)}=\mu_{t}\cdot P_{\text{recept}}^{(i,j)}\cdot\eta_{d}.$ (22)
In addition to signal photons, the receiver also detects background photons that originate from ambient light sources, thermal noise, or detector dark counts. These events are independent of the entangled source and are modeled as a separate Poisson process with mean arrival rate $\mu_{b}$. Because background photons have random polarization states, each such photon has a probability of $1/2$ to appear in either polarization-sensitive SPAD channel. Therefore, the total number of photons impinging on the receiver in slot $k$ is the sum of signal and background counts
$N_{u}[k]=N_{s}[k]+N_{b}[k],$ (23)
where $N_{s}[k]\sim\text{Poisson}(\mu_{s}^{(i,j)})$ and $N_{b}[k]\sim\text{Poisson}(\mu_{b})$ are independent.
At the user receiver, entangled photons are captured and registered over the same quantum slot structure $\mathcal{S}=\{1,2,\dots,N_{s}\}$ as defined at the reference side. The receiver constructs a binary detection sequence $\mathbf{d}_{u}=[d_{u}[1],d_{u}[2],\dots,d_{u}[N_{s}]]$, where each element $d_{u}[k]\in\{0,1,\texttt{fail}\}$ reflects the outcome of photon detection and polarization measurement during slot $k$. Let us define the detection bit $d_{u}[k]$ for slot $k$ as follows:
  • $d_{u}[k]=1$ if exactly one signal photon is detected and its polarization is correctly resolved.
  • $d_{u}[k]=0$ if a signal photon is detected but the measured polarization is incorrect (due to overlap with background).
  • $d_{u}[k]=\texttt{fail}$ in all other cases: either no photon was detected or multiple photons caused measurement ambiguity.
A slot is most likely to result in a fail status due to the absence of any detected photons (i.e., $N_{u}[k]=0$). However, even if a signal photon is present, a fail can also occur if:
  1. 1.
    A background photon arrives simultaneously with a signal photon ($N_{s}[k]=1$, $N_{b}[k]\geq 1$)
  2. 2.
    The background photon’s polarization differs from that of the signal, leading to a contradiction in polarization detection.
In quantum slots where $d_{u}[k]\in\{0,1\}$ (i.e., photon detection is successful and not marked as fail), the user receiver records the photon arrival time $t_{u}[k]$. This timestamp captures three main factors:
  • The photon generation time $t_{g}[k]$ at the transmitter (see (19))
  • A cumulative synchronization deviation $\tau_{\text{sync}}$, which includes both the physical propagation delay and any initial offset due to system synchronization mismatch
  • The SPAD detector’s timing jitter $n_{u}[k]\sim\mathcal{N}(0,\sigma_{d}^{2})$, modeled as Gaussian noise
Accordingly, for each valid detection slot $k$, the timestamp recorded at the user is given by
$t_{u}[k]=t_{g}[k]+\tau_{\text{sync}}+n_{u}[k],\quad n_{u}[k]\sim\mathcal{N}(0,\sigma_{d}^{2}).$ (24)
Here, $t_{g}[k]$ is itself a Gaussian-distributed variable centered around the pump pulse time $t_{p}[k]$ (as described in (19)). Thus, the total timing uncertainty at the receiver combines the intrinsic generation jitter $\sigma_{t}$ and the detector-side jitter $\sigma_{d}$, resulting in the following distribution
$t_{u}[k]\sim\mathcal{N}(t_{p}[k]+\tau_{\text{sync}},\sigma_{t}^{2}+\sigma_{d}^{2}).$ (25)
Finally, the complete user-side timestamp sequence is represented as
$\mathbf{t}_{u}=[t_{u}[1],t_{u}[2],\dots,t_{u}[N_{s}]],$ (26)
where each element $t_{u}[k]$ is defined only for slots where $d_{u}[k]\in\{0,1\}$, and marked as fail otherwise.
Algorithm 1 Quantum Synchronization via Sparse Bit Matching
1:Input: User detection vector $\mathbf{d}_{u}$, user timestamps $\mathbf{t}_{u}$, reference detection vector $\mathbf{d}_{r}$, reference timestamps $\mathbf{t}_{r}$
2:Output: Estimated synchronization offset $\hat{\tau}_{\text{sync}}$
3:Extract index set $\mathbf{n}_{u}=\{i_{1},\dots,i_{x}\}$ such that $d_{u}[i_{j}]\in\{0,1\}$
4:Extract bit values $\mathbf{v}_{u}=[d_{u}[i_{1}],\dots,d_{u}[i_{x}]]$
5:for each shift $\ell$ in candidate set $\mathcal{L}$ do
6:  $S[\ell]\leftarrow 0$
7:  for $j=1$ to $x$ do
8:   if $d_{r}[i_{j}-\ell]\in\{0,1\}$ and $d_{r}[i_{j}-\ell]=d_{u}[i_{j}]$ then
9:     $S[\ell]\leftarrow S[\ell]+1$
10:   end if
11:  end for
12:end for
[…]
14:Construct matching index set $\mathcal{J}_{\text{match}}=\{j:d_{r}[i_{j}-\ell^{\star}]=d_{u}[i_{j}]\}$
[…]
16:return $\hat{\tau}_{\text{sync}}$
IV-C Synchronization Recovery at the User
To establish a shared temporal reference between the transmitter and the user receiver, the system utilizes a classical auxiliary synchronization channel in addition to the quantum photon link. This classical link, implemented using technologies such as vertical-cavity surface-emitting laser (VCSEL)-based narrow optical beams, is made feasible by the structured beam-steering setup already present at the transmitter [19]. Through this channel, the transmitter broadcasts two sequences aligned with the quantum time slots, i.e., the binary detection sequence $\mathbf{d}_{r}=[d_{r}[1],\dots,d_{r}[N_{s}]]$ and the associated timestamp sequence $\mathbf{t}_{r}=[t_{r}[1],\dots,t_{r}[N_{s}]]$, which represent detection outcomes and timing information of reference photons, respectively.
At the user side, upon receiving photons through the quantum aperture, the receiver builds its own detection sequence $\mathbf{d}_{u}=[d_{u}[1],\dots,d_{u}[N_{s}]]$ and timestamp vector $\mathbf{t}_{u}=[t_{u}[1],\dots,t_{u}[N_{s}]]$, as previously described. However, due to the unknown synchronization offset $\tau_{\text{sync}}$ between local and transmitter clocks, the timestamps in $\mathbf{t}_{u}$ are shifted relative to those in $\mathbf{t}_{r}$. This alignment is further complicated by the fact that a significant portion of transmitted photons result in fail states at the receiver, primarily due to the combined effects of source-side timing jitter, SPAD detector noise, and interference from background photons. As a result, only a small subset of quantum slots provides reliable and temporally matched detection events, making accurate synchronization estimation more challenging under practical conditions.
Due to the random nature of photon detection, the binary sequence $\mathbf{d}_{u}=[d_{u}[1],\dots,d_{u}[N_{s}]]$ constructed at the user side includes a large number of fail entries, corresponding to undetected or ambiguous photon events. In contrast, the reference sequence $\mathbf{d}_{r}=[d_{r}[1],\dots,d_{r}[N_{s}]]$ generated at the transmitter is structurally aligned with the quantum slot framework, although it may also contain fail entries due to multi-photon emissions or null generation events. Nevertheless, the fail rate in $\mathbf{d}_{r}$ is typically much lower than that observed in the user-side sequence $\mathbf{d}_{u}$. The primary challenge is that the valid bits in $\mathbf{d}_{u}$ represent only a sparse and potentially misaligned sample of the original transmission pattern. Additionally, some valid detections at the user may be corrupted due to background-induced bit errors.
Considering the above-mentioned challenges, and to enable robust synchronization, we extract two auxiliary vectors from $\mathbf{d}_{u}$ as follows
  • A valid data vector $\mathbf{v}_{u}=[d_{u}[i_{1}],d_{u}[i_{2}],\dots,d_{u}[i_{x}]]$, consisting of all elements where $d_{u}[i_{j}]\in\{0,1\}$,
  • A corresponding index vector $\mathbf{n}_{u}=[i_{1},i_{2},\dots,i_{x}]$, recording the slot positions of these valid entries.
We then perform pattern matching between the observed vector $\mathbf{v}_{u}$ and shifted versions of the transmitter-side reference vector $\mathbf{d}_{r}$, evaluating how well $\mathbf{v}_{u}$ aligns with subsequences of $\mathbf{d}_{r}$. For each candidate integer shift $\ell\in\mathcal{L}$, we compute a matching score as
$S[\ell]=\sum_{j=1}^{x}\delta\left(d_{u}[i_{j}],d_{r}[i_{j}-\ell]\right),$ (27)
where $\delta(a,b)=1$ if $a=b$, and 0 otherwise. The optimal alignment offset is then estimated as
$\ell^{\star}=\arg\max_{\ell\in\mathcal{L}}S[\ell],$ (28)
which identifies the temporal shift that yields the highest agreement between the user and reference detection sequences, ignoring fail entries.
Once the optimal shift $\ell^{\star}$ is identified, the user constructs a refined index set $\mathcal{J}_{\text{match}}$ by filtering those valid entries for which both transmitter and receiver have defined and matching binary values
$\mathcal{J}_{\text{match}}=$ (29)
$\left\{j\in\{1,\dots,x\}\;:\;d_{r}[i_{j}-\ell^{\star}]\in\{0,1\},\;d_{r}[i_{j}-\ell^{\star}]=d_{u}[i_{j}]\right\}.$
This filtered subset ensures that the timestamp estimation only uses reliably matched positions, where both the quantum event occurred and the polarization was correctly inferred at the receiver. The final synchronization offset is then estimated by computing the average temporal difference over the selected indices, i.e.,
$\hat{\tau}_{\text{sync}}=\frac{1}{|\mathcal{J}_{\text{match}}|}\sum_{j\in\mathcal{J}_{\text{match}}}\left(t_{u}[i_{j}]-t_{r}[i_{j}-\ell^{\star}]\right).$ (30)
This two-stage matching and refinement process enables robust synchronization even in the presence of asymmetric detection failures, background-induced bit flips, and missing reference values. It ensures that only high-confidence matchings contribute to the final clock alignment.
Figure 2: Effect of synchronization sequence duration $T_{\text{seq}}$ and grid configuration on synchronization performance. (a) Probability of synchronization failure and (b) Average synchronization error versus $T_{\text{seq}}$ under different spatial grid sizes. A failure occurs when fewer than two valid detection bits are registered within the allotted duration, making offset estimation infeasible.
Based on the statistical detection and timing models defined in (24)–(25) and the matching score formulation in (27)–(28), the proposed synchronization strategy is summarized in Algorithm 1. This procedure leverages a two-stage approach. First, it estimates the most likely alignment shift, $\ell^{\star}$, between the sparse user-side detection vector $\mathbf{d}_{u}$ and the reference pattern $\mathbf{d}_{r}$ using discrete correlation. Then, it constructs a refined set of matching slots $\mathcal{J}_{\text{match}}$ that excludes ambiguous or non-informative events. The synchronization offset $\hat{\tau}_{\text{sync}}$ is subsequently computed via averaging over timestamp differences as described in (30). This method allows reliable clock recovery even under heavy detection sparsity and measurement noise, and forms the basis for slot-level temporal alignment required for entanglement correlation and key distribution in quantum communication scenarios.
V Simulations
To assess the synchronization performance under realistic quantum indoor environments, we conducted Monte-Carlo simulations based on the system model described in Sections IIIV. The room was configured as a rectangular indoor space with dimensions $L=6\,\text{m}$, $W=6\,\text{m}$, and $H=3\,\text{m}$. The quantum optical coverage was set at height $h_{\text{cov}}=1\,\text{m}$ above the floor, and the grid-based beam steering was evaluated for multiple spatial configurations with $N_{g}\in\{25,100,225\}$, corresponding to $5\times 5$, $10\times 10$, and $15\times 15$ grid layouts. Each grid region was targeted by a distinct optical beam, and photon propagation followed the diffraction and spatial uncertainty models described earlier.
The synchronization sequence spanned a total duration $T_{\text{seq}}$ ranging from $1\,\mu\text{s}$ to $2\,\text{ms}$, divided into slots of duration $T_{qs}=10\,\text{ns}$. This results in a total number of time slots $N_{s}=T_{\text{seq}}/T_{qs}$, which varied between $10^{2}$ and $2\times 10^{5}$ depending on the experiment. Each slot could contain a single entangled photon pair with mean generation rate $\mu_{t}=0.5$, while the SPAD detector timing jitter was modeled as Gaussian noise with standard deviation $\sigma_{d}=200\,\text{ps}$. The SPDC source temporal jitter was modeled with standard deviation $\sigma_{t}=50\,\text{ps}$ [35]. Background photon arrivals followed a Poisson distribution with mean $\mu_{b}=5\times 10^{-5}$. The quantum detection efficiency of the SPAD was set to $\eta_{d}=0.6$, and the user receiver was modeled with a circular aperture of radius $r_{a}=0.02\,\text{m}$. The simulation parameters are summarized in Table II, where variables marked with a range were evaluated parametrically, and exact values for each scenario are specified in the respective figure captions.
TABLE II: Simulation Parameters for Quantum Synchronization System
Parameter Symbol Value / Range
Room length $L$ $6\,\text{m}$
Room width $W$ $6\,\text{m}$
Room height $H$ $3\,\text{m}$
Coverage height $h_{\text{cov}}$ $1\,\text{m}$
Number of grids $N_{g}$ $25$, $100$, $225$
Beam width per grid $w_{z}^{(i,j)}$ $\frac{W}{N_{x}}$
Receiver aperture radius $r_{a}$ $0.02\,\text{m}$
Slot duration $T_{qs}$ $10\,\text{ns}$
[…]
Number of time slots $N_{s}$ $10^{2}$$2\times 10^{5}$
Mean photon-pair rate per slot $\mu_{t}$ $0.5$
SPAD detection efficiency $\eta_{d}$ $0.6$
Mean background photons per slot $\mu_{b}$ $5\times 10^{-6}$
SPDC temporal jitter $\sigma_{t}$ $50\,\text{ps}$
SPAD timing jitter $\sigma_{d}$ $200\,\text{ps}$
Wavelength of photon $\lambda$ $1550\,\text{nm}$
Initial position error std. dev. $\sigma_{p}$ $6\,\text{cm}$
Fig. 2 evaluates the effect of synchronization sequence duration $T_{\text{seq}}$ on the quantum synchronization performance under different spatial grid configurations. Each subfigure corresponds to a distinct performance metric, with subfigure (a) showing the probability of synchronization failure, and subfigure (b) reporting the average synchronization error between the user and transmitter time references. As illustrated in Fig. 2a, the probability of synchronization failure, which is defined as the event where fewer than two valid detection bits are observed at the user within the allotted synchronization time, decreases monotonically with increasing $T_{\text{seq}}$. This phenomenon arises due to the inherent sparsity and stochastic nature of photon reception in wireless quantum systems. Given the probabilistic losses in entangled photon transmission and background interference, many quantum slots yield no valid detections. When the synchronization duration is insufficient, the receiver may fail to collect the minimum required number of valid qubits (two or more), leading to a failure in offset estimation. Fig. 2b further illustrates that the mean synchronization error, which is defined as the average deviation between the estimated synchronization offset $\hat{\tau}_{\text{sync}}$ and the true offset $\tau_{\text{sync}}$, declines significantly as $T_{\text{seq}}$ increases. A longer sequence provides more detection opportunities, enhancing the likelihood of high-confidence bit matching and robust timestamp averaging. Additionally, both metrics demonstrate that finer spatial grid configurations (e.g., $15\times 15$) lead to better synchronization outcomes compared to coarser grids (e.g., $5\times 5$). This improvement is due to enhanced beam alignment, which increases the photon reception probability $P_{\text{recept}}^{(i,j)}$ (as derived in (16)). With narrower beam footprints, users are more likely to fall within high-intensity zones, enabling valid qubit detection even under shorter synchronization durations. For instance, after $T_{\text{seq}}=1\,\text{ms}$, the average synchronization error is reduced from approximately $90\,\text{ps}$ in the $5\times 5$ case to below $9\,\text{ps}$ in the $15\times 15$ configuration.
Figure 3: Average synchronization error versus photon-pair generation rate $\mu_{t}$ for grid size $15\times 15$ and synchronization duration $T_{\text{seq}}=1\,\text{ms}$.
Another critical parameter that significantly affects the performance of wireless quantum synchronization systems is the mean photon-pair generation rate per slot, denoted by $\mu_{t}$. Fig. 3 illustrates the average synchronization error as a function of $\mu_{t}$ over a wide range from $0.1$ to $4$, under fixed environmental and timing configurations. The simulation considers a $15\times 15$ grid structure ($N_{g}=225$) and a synchronization duration of $T_{\text{seq}}=1\,\text{ms}$ per user. All other system parameters follow the nominal values specified in Table II. As evident in the results, the value of $\mu_{t}$ plays a dual and non-trivial role in synchronization accuracy. When $\mu_{t}$ is too small, the probability that the SPDC source generates a valid single photon pair in a given quantum slot decreases sharply, leading to high rates of slot failure and consequently degraded synchronization performance due to insufficient valid timestamps. Conversely, if $\mu_{t}$ becomes too large, the probability of generating two or more photon pairs in the same slot increases, which, according to the protocol defined in (17), renders the slot invalid for synchronization purposes. Such multi-pair emissions introduce ambiguity and break the one-to-one correspondence between reference and user photons. Therefore, there exists an optimal range for $\mu_{t}$ that balances the trade-off between too few photon events and too many invalid multiphoton occurrences. As shown in Fig. 3, this balance is achieved near $\mu_{t}\approx 1$, which minimizes the average synchronization error. Selecting $\mu_{t}$ carefully around this optimal region is essential for reliable and high-precision quantum synchronization under constrained photon budgets and noisy channel conditions.
VI Conclusion and Future Work
In this work, we introduced a grid-structured, entanglement-based quantum synchronization architecture for indoor wireless optical systems. The proposed framework leverages a central ceiling-mounted SPDC transmitter that steers time-energy entangled photons toward spatially targeted beams, each corresponding to a predefined spatial grid. By estimating the user’s location and aligning the optical beam accordingly, the system increases the likelihood of photon reception despite the probabilistic nature of quantum propagation and detection. A beam-local coordinate transformation was applied to analytically model the spatial photon reception probability, which plays a key role in determining synchronization reliability. To overcome sparsity and randomness in photon arrivals, we proposed a robust two-stage synchronization recovery algorithm based on sparse bit-pattern matching and timestamp alignment. This method enables accurate estimation of clock offsets even in the presence of background noise, multiphoton events, and detector jitter. Simulation results demonstrated the effectiveness of the approach under practical conditions. It was shown that increasing the synchronization sequence duration significantly reduces both synchronization failure probability and mean timing error. Furthermore, increasing grid resolution from $5\times 5$ to $15\times 15$ notably improves photon alignment and enhances synchronization accuracy, achieving sub-100 ps mean error within $1\,\text{ms}$ of allocated time. Additionally, we evaluated the effect of photon-pair generation rate $\mu_{t}$ and identified an optimal range that minimizes timing error by balancing between insufficient single-pair events and excess multiphoton emissions.
Future work will focus on extending the system to multi-transmitter environments to enhance spatial coverage, support user mobility, and improve robustness against obstructions. Additionally, a promising direction involves integrating synchronization with quantum communication, sensing, and positioning into a unified framework. Such a system could enable sub-centimeter positioning and sub-picosecond synchronization accuracy, far surpassing the capabilities of conventional classical techniques, and pave the way for secure, high-precision, quantum-enabled indoor wireless infrastructures.
Appendix A
Based on (2), we have
$\hat{\mathbf{z}}^{\prime}=\frac{\mathbf{d}^{(i,j)}}{\|\mathbf{d}^{(i,j)}\|}.$ (31)
To construct the orthonormal basis for the local coordinate system, we choose a reference vector (not parallel to $\hat{\mathbf{z}}^{\prime}$), e.g.,
$\mathbf{v}_{\text{ref}}=\begin{cases}[0,0,1],&\text{if }\hat{\mathbf{z}}^{\prime}\text{ is not close to }[0,0,1]\\ [1,0,0],&\text{otherwise}\end{cases}$ (32)
Then define:
$\hat{\mathbf{x}}^{\prime}$ $=\frac{\mathbf{v}_{\text{ref}}\times\hat{\mathbf{z}}^{\prime}}{\|\mathbf{v}_{\text{ref}}\times\hat{\mathbf{z}}^{\prime}\|},$ (33)
$\hat{\mathbf{y}}^{\prime}$ $=\hat{\mathbf{z}}^{\prime}\times\hat{\mathbf{x}}^{\prime}.$ (34)
The resulting rotation matrix from global $(x,y,z)$ to beam-aligned $(x^{\prime},y^{\prime},z^{\prime})$ coordinates is:
$\mathbf{R}^{(i,j)}=\left[\hat{\mathbf{x}}^{\prime\top}\;\hat{\mathbf{y}}^{\prime\top}\;\hat{\mathbf{z}}^{\prime\top}\right]^{\top}$ (35)
Any point $\mathbf{p}$ in the global coordinate system is then transformed into the beam-aligned frame as:
$\mathbf{p}^{\prime}=\mathbf{R}^{(i,j)}\left(\mathbf{p}-\mathbf{p}_{\text{tx}}\right)$ (36)